### v3

# Competition between private and expressed opinions in binary choice: the $\alpha$-EPO $q$-voter model

Abstract
People often express opinions that differ from their privately held views, a phenomenon known in economy as preference falsification. Expressed-private opinion (EPO) models capture this by assigning each agent two dynamical variables: a private (internal) and an expressed (external) opinion. Within the nonlinear $q$-voter model, two EPO variants have been studied so far: with and without self-anticonformity. In both formulations, agents update private and expressed binary opinions, one after another and at the same rate, which has led to two update schemes studied previously: AT (act then think), in which an agent first updates its expressed and then its private opinion, and TA (think then act), in which the order is reversed. To eliminate this ad hoc distinction and quantify the interplay between private and expressed opinions, we introduce the $\alpha$-EPO $q$-voter model with asynchronous updating – in each elementary step, an agent updates its private opinion with probability $\alpha$ or its expressed opinion with complementary probability $1-\alpha$. We derive mean-field theory and, for the first time for EPO $q$-voter dynamics, a pair approximation, and validate them with Monte Carlo simulations on artificial and real organizational networks. Comparing the two model variants, we show that the collective outcome controlled by $\alpha$ strongly depends on self-anticonformity: with self-anticonformity the results are robust to $\alpha$, whereas without it $\alpha$ shifts the agreement–disagreement threshold and, for $q=3$, changes the phase transition from continuous to discontinuous. Based on these results, we briefly discuss implications for real societies that may be relevant to managers and policymakers.
I Introduction
It is well known that people do not always express their true preferences, a phenomenon known as preference falsification, because they may fear social disapproval, sanctions, or negative consequences arising from openly stating views that deviate from perceived social norms [20]. A natural modeling approach to such situations is to distinguish between an agent’s private and expressed states. This idea underlies the class of expressed-private opinion (EPO) models [4]; for a short and recent review, see [16].
Here, we focus on a specific aspect of these models: the effect of unequal update rates for private and expressed opinions in a binary choice framework, which is relevant to surprisingly complex problems [35]. We study this effect within the nonlinear $q$-voter model, one of the most popular opinion-dynamics models introduced in physics [3, 30, 28, 11]. There are various approaches to modeling the interplay between private and expressed (often called public) opinions. Some models assume that private opinions do not change over time [26, 8, 25, 24]. By contrast, others allow both behavior and attitude to change at the same rate [1, 37, 21, 12, 4, 31]. Finally, there are models in which the update rates for expressed and private opinions differ. An interesting example is the model proposed by Piras et al. [33], in which the agents are generally more likely to change their behavior than their attitudes. However, when agents experience strong cognitive dissonance [6], whether they change their behavior or opinion depends on how often they have altered their opinions in the past. Another example, much closer to what we propose in this paper, is the concealed voter model [9], in which agents change their expressed opinion by copying the opinion of a randomly selected neighbor (with probability $c$) or externalizing their internal opinion (with probability $e$), while the private opinion is updated by internalizing the expressed one (with probability $i$).
Despite the fact that $q$-voter models [3, 30, 28, 11] typically follow an asynchronous updating scheme – at each elementary time step a single voter updates its opinion – the two $q$-voter-based EPO models [13, 16] assume that private and expressed opinions are updated in a quasi-synchronous manner. In this formulation, at each step a single voter updates both its private and public opinion in a specified order. Within this updating scheme, it has been shown that the stationary average private and expressed opinions (i.e., the collective-level aggregates) are independent of whether agents update their private opinion before their expressed opinion, or vice versa [13, 16]. It remains unclear, however, whether introducing rate asymmetry between private and expressed opinions alters the collective dynamics. We therefore extend these models by introducing an asynchronous update scheme, in which only one layer is updated at each step: private opinions are updated with probability $\alpha$, and expressed opinions with the complementary probability $1-\alpha$. We refer to these extensions as $\alpha$-EPO models. We study both variants considered in  [13, 16] and analyze the resulting dynamics on synthetic and empirical networks using mean-field and pair approximation, as well as Monte Carlo simulations.
Studying both variants in parallel is particularly interesting because, as recently suggested, the older version of the model can be interpreted as incorporating a form of self-anticonformity [13], whereas the newer version does not include this mechanism [16]. Furthermore, Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16] considered their models only on a complete graph, corresponding to a fully connected society in which everyone knows everyone. For large social systems, this assumption is clearly unrealistic. Here, we move beyond the complete graph by studying $\alpha$-EPO $q$-voter models on networks using Monte Carlo (MC) simulations and pair approximation (PA) analysis.
II The model
We consider a system of $N$ agents, indexed by $i=1,\ldots,N$, each described by two binary variables: $S_{i}(t)\in\{-1,+1\}$, the expressed opinion of agent $i$ at time $t$ (visible to other agents), and $\sigma_{i}(t)\in\{-1,+1\}$, the private opinion of agent $i$ at time $t$ (not visible to others). The binary variables represent a choice between two alternatives, such as yes or no; for or against; adopt or not adopt, etc. Agents occupy the nodes of an undirected, unweighted network, whose topology specifies the interaction patterns within the system. For each agent $i$, let $\mathcal{N}(i)$ denote the set of its neighbors and let $k_{i}=|\mathcal{N}(i)|$ be the degree of node $i$; thus, agent $i$ can interact only with agents in $\mathcal{N}(i)$. In conformity events, the influence group (the $q$-panel) is formed by selecting uniformly at random, without replacement, $q$ neighbors of agent $i$, i.e.,a subset $\mathcal{Q}\subset\mathcal{N}(i)\;\wedge\;|\mathcal{Q}|=q$. In all analytical calculations and algorithmic description of the model we assume that every node in the network has at least $q$ neighbors.
Agents can update their opinions at both levels (expressed and private) due either to social influence from their neighbors or to the need to reduce internal dissonance, which arises when private and expressed opinions are in conflict. The updating rules for public and private opinions differ; however, at both levels we consider two mechanisms: independence and conformity. Independence at the expressed level means that an agent simply verbalizes its private opinion, i.e., its expressed opinion is set equal to the private one, thereby reducing the agent’s internal dissonance. In contrast, independence at the private level corresponds to stochastic noise, as in many other versions of the nonlinear voter model [2, 17, 18, 32, 36]. Specifically, we use the formulation proposed in [30]: under independence, an agent flips its private opinion to the opposite one with probability $1/2$.
Conformity at both levels is driven by social pressure exerted by other agents. In our models, the source of this social influence is a group of $q$ randomly selected neighbors, called the $q$-panel. At the expressed level, conformity comprises two scenarios, depending on the agent’s initial state. If an agent is initially in harmony (i.e., its private and expressed opinions coincide), it is less susceptible to influence. Therefore, it adjusts its expressed opinion to match the group only if the $q$-panel is unanimous, that is, if all $q$ selected neighbors share the same expressed opinion. This type of behavior is called compliance [29]. By contrast, an agent who is in dissonance and thus experiences internal tension is more willing to change. In this case, it is enough that at least one member of the $q$-panel publicly expresses an opinion consistent with the agent’s private beliefs to encourage the agent to speak up, align its expressed opinion with its private one, and thereby relieve dissonance. This type of response to social influence is referred to as disinhibitory contagion [29], a phenomenon in which breaking the norm by even a single individual incites others to act according to their true intentions.
Conformity at the private level also requires a source of influence. Here, a key difference emerges between the models proposed by Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16], and this distinction carries over to the two variants considered in this paper. In the first case, following Jędrzejewski et al. [13], an agent conforms to a unanimous $q$-panel regardless of its expressed opinion. Consequently, the agent may change its private opinion to match the group while still expressing the opposite view. We refer to this as self-anticonformity, since the agent, at the expressed level, disagrees not only with the group but also with its own private opinion. In the second variant, following Kamińska and Sznajd-Weron [16], this scenario is prohibited. In other words, an agent conforms privately only if the $q$-panel is unanimous and its opinion matches the agent’s expressed one. This also implies that conformity on the private level cannot lead to dissonance.
In this paper, we introduce an extension that applies to both variants of the model described above, allowing for a more flexible relationship between updates at the expressed and private levels. The proposed extension captures the fact that a change in private opinion is not necessarily followed by a change in behavior, and vice versa: behavior may change without altering beliefs [19]. Both Jędrzejewski et al. [13], Kamińska and Sznajd-Weron [16] assumed that during a single update, both expressed and private opinions are updated. They considered two update schemes: “act then think” (AT), in which the expressed opinion is updated first and the private one second, and “think then act” (TA), in which the private opinion is updated first and the expressed one second. Here, we assume that with probability $\alpha$ an agent updates its private opinion, and with complementary probability $1-\alpha$ it updates its expressed opinion. The algorithm for an elementary update of the $\alpha$-EPO model, taking time $\Delta t$, is described by Algorithm 1 and illustrated in Figs. 1 and 2.
Algorithm 1 Elementary opinion update
1:Draw target agent $i\sim U\{1,\ldots,N\}$
2:Draw $r_{\alpha}\sim U(0,1)$
3:if $r_{\alpha}<\alpha$ then $\triangleright$ Update private opinion $\sigma_{i}$
4:  Draw $r_{p}\sim U(0,1)$
5:  if $r_{p}<p$ then $\triangleright$ Independence
6:   $\sigma_{i}(t+\Delta t)\leftarrow(-1)^{x}\text{, where }x\sim\text{Bernoulli}\left(\frac{1}{2}\right)$
[…]
8:   Draw random set $\mathcal{Q}\subset\mathcal{N}(i)$ such that $|\mathcal{Q}|=q$
9:   if model with self-anticonformity then
10:     if $\forall j\in\mathcal{Q},\;S_{j}(t)=-\sigma_{i}(t)$ then
11:      $\sigma_{i}(t+\Delta t)\leftarrow-\sigma_{i}(t)$
12:     else
13:      $\sigma_{i}(t+\Delta t)\leftarrow\sigma_{i}(t)$
14:     end if
15:   else if model without self-anticonformity then
16:     if $\forall j\in\mathcal{Q},\;S_{j}(t)=S_{i}(t)$ then
17:      $\sigma_{i}(t+\Delta t)\leftarrow S_{i}(t)$
18:     else
19:      $\sigma_{i}(t+\Delta t)\leftarrow\sigma_{i}(t)$
20:     end if
21:   end if
22:  end if
[…]
24:  Draw $r_{p}\sim U(0,1)$
25:  if $r_{p}<p$ then $\triangleright$ Independence
26:   $S_{i}(t+\Delta t)\leftarrow\sigma_{i}(t)$
[…]
28:   Draw random set $\mathcal{Q}\subset\mathcal{N}(i)$ such that $|\mathcal{Q}|=q$
29:   if $S_{i}(t)\neq\sigma_{i}(t)$ then $\triangleright$ Disinhibitory contagion
30:     if $\exists j\in\mathcal{Q}\text{ such that }S_{j}(t)=\sigma_{i}(t)$ then
31:      $S_{i}(t+\Delta t)\leftarrow\sigma_{i}(t)$
32:     else
33:      $S_{i}(t+\Delta t)\leftarrow S_{i}(t)$
34:     end if
[…]
36:     if $\forall j\in\mathcal{Q},\;S_{j}(t)=-S_{i}(t)$ then
37:      $S_{i}(t+\Delta t)\leftarrow-S_{i}(t)$
38:     else
39:      $S_{i}(t+\Delta t)\leftarrow S_{i}(t)$
40:     end if
41:   end if
42:  end if
43:end if
[…]
Figure 1: Visualization of a single update in the models with and without self-anticonformity (see the description directly above the bottom row in the left panel), for the case in which the focal agent is in harmony before the update.
Figure 2: Visualization of a single update in the models with and without self-anticonformity (see the description directly above the bottom row in the left panel), for the case in which the focal agent is in dissonance before the update.
Let us, following [13, 16], use $\uparrow$ to denote the opinion +1+1 and $\downarrow$ to denote the opinion 1-1. To describe the state of a single agent at time $t$, we use the ordered pair $(S(t),\sigma(t))$, where the first element is the expressed opinion, followed by the private one. Accordingly, the set of possible agent states is $X=(S,\sigma)\in\{\uparrow\uparrow,\uparrow\downarrow,\downarrow\uparrow,\downarrow\downarrow\}$. The macroscopic state of the system at time $t$ can be described by the fractions of agents in each state, denoted by $c_{X}(t)$, defined as
$c_{X}(t)=\frac{N_{X}(t)}{N},$ (1)
where $N_{X}(t)$ denotes the number of agents in state $X$ at time $t$, and $N$ is the total number of agents in the system. Since the total number of agents in the system is fixed, we have
$c_{\uparrow\uparrow}(t)+c_{\uparrow\downarrow}(t)+c_{\downarrow\uparrow}(t)+c_{\downarrow\downarrow}(t)=1.$ (2)
To enable comparison with Jędrzejewski et al. [13] and Kamińska and Sznajd-Weron [16] instead of analyzing the concentrations of agents in all possible states separately, we consider the following quantities:
$c_{S}(t)=c_{\uparrow\uparrow}(t)+c_{\uparrow\downarrow}(t),$ (3)
being the fraction of agents with positive expressed opinion,
$c_{\sigma}(t)=c_{\uparrow\uparrow}(t)+c_{\downarrow\uparrow}(t),$ (4)
being the fraction of agents with positive private opinion, and
$d(t)=c_{\uparrow\downarrow}(t)+c_{\downarrow\uparrow}(t),$ (5)
being the fraction of agents in dissonance.
III Methods
III.1 Monte Carlo simulations
We use Monte Carlo simulations both to validate our analytical calculations for complete and random graphs, as well as to extend the analysis to other topologies, including the empirical networks. To obtain smooth trajectories comparable with the mean-field and pair approximations (which effectively correspond to the thermodynamic limit), we simulate relatively large systems with $N=10^{4}$ and $N=10^{5}$ nodes. We also conduct MC simulations on empirical networks from organizations [7]. In the latter case, some nodes may have degree lower than $q$. In such cases we assume that the $q$-panel consists of all neighbors, i.e., for each agent $i$ we take $q_{i}=\min(k_{i},q)$.
In our MC simulations, one MC step (MCS) consists of $N$ elementary updates performed according to the algorithm described in Section II and illustrated in Figs. 1 and 2. We assume that agents are initially in harmony, i.e., $\forall_{i};S_{i}(0)=\sigma_{i}(0)$. We use two types of initial conditions: ordered, $c_{S}(0)=c_{\sigma}(0)=1$, and random, where $c_{S}(0)=c_{\sigma}(0)=1/2$. After initialization, we equilibrate the system for 3000 MCS, and then perform time averaging over the subsequent 2000 MCS. The simulation codes are publicly available at github.com/BarbaraKaminska/EPO-q-voter .
III.2 Mean field approximation
As stated above, one Monte Carlo step consists of $N$ elementary updates, hence $N\Delta t=1$, which implies $\Delta t=1/N$. During a single elementary update, the fraction of agents in a given state can increase by $1/N$, decrease by $1/N$, or remain unchanged. We define the probabilities of these changes as
$\gamma_{X}^{+}(t)$ $=P(c_{X}(t+1/N)=c_{X}(t)+1/N),$ (6)
$\gamma_{X}^{-}(t)$ $=P(c_{X}(t+1/N)=c_{X}(t)-1/N),$ (7)
$\gamma_{X}^{0}(t)$ $=P(c_{X}(t+1/N)=c_{X}(t)).$ (8)
This yields a system of four equations. Below, we write it in explicit form. To simplify notation, in what follows we omit the explicit time argument $t$ whenever no confusion may arise; in particular, we write $c_{X}$ instead of $c_{X}(t)$ (and analogously for $\gamma_{X}^{\pm}$, $S_{i}$, and $\sigma_{i}$). Time dependence will be stated explicitly only when needed.
$\frac{dc_{\uparrow\uparrow}}{dt}$ $=c_{\uparrow\downarrow}\alpha\left[p/2+(1-p)c_{S}^{q}\right]+c_{\downarrow\uparrow}(1-\alpha)\left[1-(1-p)(1-c_{S})^{q}\right]$
$-c_{\uparrow\uparrow}\left\{1-\alpha\left[1-p/2\,\boxed{-(1-p)(1-c_{S})^{q}}\right]-(1-\alpha)\left[1-(1-p)(1-c_{S})^{q}\right]\right\},$ (9)
$\frac{dc_{\uparrow\downarrow}}{dt}$ $=c_{\uparrow\uparrow}\alpha\left[p/2\,\boxed{+(1-p)(1-c_{S})^{q}}\right]+c_{\downarrow\downarrow}(1-\alpha)(1-p)c_{S}^{q}-c_{\uparrow\downarrow}\left\{1-\alpha\left[1-p/2-(1-p)c_{S}^{q}\right]-(1-\alpha)(1-p)c_{S}^{q}\right\},$ (10)
$\frac{dc_{\downarrow\uparrow}}{dt}$ $=c_{\uparrow\uparrow}(1-\alpha)(1-p)(1-c_{S})^{q}+c_{\downarrow\downarrow}\alpha\left[p/2\,\boxed{+(1-p)c_{S}^{q}}\right]$
$-c_{\downarrow\uparrow}\left\{1-\alpha\left[1-p/2-(1-p)(1-c_{S})^{q}\right]-(1-\alpha)(1-p)(1-c_{S})^{q}\right\},$ (11)
$\frac{dc_{\downarrow\downarrow}}{dt}$ $=c_{\uparrow\downarrow}(1-\alpha)\left[1-(1-p)c_{S}^{q}\right]+c_{\downarrow\uparrow}\alpha[p/2+(1-p)(1-c_{S})^{q}]$
$-c_{\downarrow\downarrow}\left\{1-\alpha\left[1-p/2\,\boxed{-(1-p)c_{S}^{q}}\right]-(1-\alpha)\left[1-(1-p)c_{S}^{q}\right]\right\}.$ (12)
In Eqs.(9)-(12), the terms enclosed in frames indicate contributions that are absent in the variant without self-anticonformity. Consequently, the equations for the model without self-anticonformity are obtained by simply dropping the framed terms. These contributions correspond to updates in which private-level conformity leads to dissonance.
The stationary state, that we are interested in, is defined by
$\frac{dc_{X}}{dt}=0$ (13)
for all $X$. Although the resulting system can be written in a fairly concise form, it must be solved numerically.
III.2.1 Model with self-anticonformity
Below we simplify the mean-field equations for the stationary states for the model with self-anticonformity.
$c_{\uparrow\uparrow}$ $=c_{\sigma}(1-\alpha)\left[1-(1-p)(1-c_{S})^{q}\right]+c_{\uparrow\uparrow}\alpha\left[1-p/2\,-(1-p)(1-c_{S})^{q}\right]+c_{\uparrow\downarrow}\alpha\left[p/2+(1-p)c_{S}^{q}\right],$ (14)
$c_{\uparrow\downarrow}$ $=(1-c_{\sigma})(1-\alpha)(1-p)c_{S}^{q}+c_{\uparrow\uparrow}\alpha\left[p/2+(1-p)(1-c_{S})^{q}\right]+c_{\uparrow\downarrow}\alpha\left[1-p/2-(1-p)c_{S}^{q}\right],$ (15)
$c_{\downarrow\uparrow}$ $=c_{\sigma}(1-\alpha)(1-p)(1-c_{S})^{q}+c_{\downarrow\uparrow}\alpha\left[1-p/2-(1-p)(1-c_{S})^{q}\right]+c_{\downarrow\downarrow}\alpha\left[p/2\,+(1-p)c_{S}^{q}\right],$ (16)
$c_{\downarrow\downarrow}$ $=(1-c_{\sigma})(1-\alpha)\left[1-(1-p)c_{S}^{q}\right]+c_{\downarrow\uparrow}\alpha[p/2+(1-p)(1-c_{S})^{q}]+c_{\downarrow\downarrow}\alpha\left[1-p/2\,-(1-p)c_{S}^{q}\right].$ (17)
We can proceed analogously to Jędrzejewski et al. [13], that is we add (14) and (15), then (14) and (16). Then, using Eqs. (3) and (4) we finally obtain
$c_{S}(1-\alpha)$ $=c_{\sigma}(1-\alpha)\left[1-(1-p)\left[(1-c_{S})^{q}+c_{S}^{q}\right]\right]$
$+(1-\alpha)(1-p)c_{S}^{q},$ (18)
$c_{\sigma}\alpha$ $=c_{\sigma}\alpha\left[1-p/2-(1-p)(1-c_{S})^{q}\right]$
$+(1-c_{\sigma})\alpha\left[p/2+(1-p)c_{S}^{q}\right].$ (19)
Hence for $\alpha\neq 0$ and $\alpha\neq 1$ we obtain exactly the same system of equations as Jędrzejewski et al. [13], namely
$c_{S}$ $=c_{\sigma}\left[1-(1-p)\left[(1-c_{S})^{q}+c_{S}^{q}\right]\right]+(1-p)c_{S}^{q},$ (20)
$c_{\sigma}$ $=\frac{p/2+(1-p)c_{S}^{q}}{p+(1-p)\left[(1-c_{S})^{q}+c_{S}^{q}\right]}.$ (21)
This means that, for the model with self-anticonformity, the parameter $\alpha$ does not affect the stationary values of the average private and expressed opinions as long as $\alpha\in(0,1)$.
Let us notice that for $\alpha=0$, Eq. (19) becomes an identity, and Eq. (18) is the same as (20). Recalling that for $\alpha=0$, agents update only their expressed opinion, the aforementioned observations mean that the fraction of agents with a positive expressed opinion depends on the initial fraction of agents with a positive private opinion. This dependence can be derived by solving eq. (20) for $c_{S}$. Analogously, Eq. (18) becomes an identity equation for $\alpha=1$. In this case, agents update only their private opinions, and we can solve Eq. (21) for $c_{\sigma}$ by plugging in the independence probability $p$ and the initial fraction of public opinion $c_{S}$.
III.2.2 Model without self-anticonformity
As stated above, for the model without self-anticonformity, the system is reduced analogously, we simply remove the framed terms, which yields:
$c_{\uparrow\uparrow}$ $=c_{\sigma}\left[1-(1-p)(1-c_{S})^{q}\right](1-\alpha)+c_{\uparrow\uparrow}\alpha\left[1-p/2\right]+c_{\uparrow\downarrow}\alpha\left[p/2+(1-p)c_{S}^{q}\right],$ (22)
$c_{\uparrow\downarrow}$ $=(1-c_{\sigma})(1-\alpha)(1-p)c_{S}^{q}+c_{\uparrow\uparrow}\alpha p/2+c_{\uparrow\downarrow}\alpha\left[1-p/2-(1-p)c_{S}^{q}\right],$ (23)
$c_{\downarrow\uparrow}$ $=c_{\sigma}(1-\alpha)(1-p)(1-c_{S})^{q}+c_{\downarrow\uparrow}\alpha\left[1-p/2-(1-p)(1-c_{S})^{q}\right]+c_{\downarrow\downarrow}\alpha p/2,$ (24)
$c_{\downarrow\downarrow}$ $=(1-c_{\sigma})(1-\alpha)\left[1-(1-p)c_{S}^{q}\right]+c_{\downarrow\uparrow}\alpha\left[p/2+(1-p)(1-c_{S})^{q}\right]+c_{\downarrow\downarrow}\alpha\left[1-p/2\right].$ (25)
By adding (22) and (23) we obtain
$c_{S}(1-\alpha)$ $=c_{\sigma}(1-\alpha)\left[1-(1-p)\left[(1-c_{S})^{q}+c_{S}^{q}\right]\right]$
$+(1-\alpha)(1-p)c_{S}^{q}.$ (26)
It is worth noting that for $\alpha\neq 1$ this simplifies to Eq. (20). However, adding Eqs. (22) and (24) does not lead to such a simple expression: the result cannot be written solely as a function of $c_{S}$, $c_{\sigma}$, and the model parameters $q$, $p$, and $\alpha$.
Again, let us briefly consider the extreme cases of $\alpha$ equal to 0 and 1. For $\alpha=0$, the result is the same as for the model with self-anticonformity, as in this scenario agents update only their expressed opinion, which evolves according to the same algorithm in both variants of the model. When $\alpha=1$, Eq. (26) becomes an identity, and consequently we cannot derive simple dependence of $c_{\sigma}$ on $c_{S}$. For $\alpha=1$, Eqs. (22)-(25) include distinct terms containing $c_{\uparrow\downarrow}$ and $c_{\downarrow\uparrow}$, which cannot be reduced to depend solely on $c_{S}$ and $c_{\sigma}$.
III.3 Pair approximation
Pair approximation is an analytical approach that relaxes the assumption of a completely homogeneous agent system by accounting for dynamical correlations between pairs of agents in different states. It has been successfully applied to many versions of the $q$-voter model, in which agents are described by a single opinion [15, 11, 36, 14, 34], but not to the EPO $q$-voter models. Within this framework, the evolution equations for the state concentrations in the model with self-anticonformity are given by the following system (for the model without self-anticonformity, the framed terms are omitted):
$\frac{dc_{\uparrow\uparrow}}{dt}$ $=(1-\alpha)\left[pc_{\downarrow\uparrow}+(1-p)\left\{c_{\downarrow\uparrow}\left(1-\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right)-c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right\}\right]$
$+\alpha\left[\frac{p}{2}(c_{\uparrow\downarrow}-c_{\uparrow\uparrow})+(1-p)\left\{c_{\uparrow\downarrow}\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}\boxed{-c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}}\right\}\right]$
$\frac{dc_{\uparrow\downarrow}}{dt}$ $=(1-\alpha)\left[-pc_{\uparrow\downarrow}+(1-p)\left\{c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}-c_{\uparrow\downarrow}\left(1-\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}\right)\right\}\right]$
$+\alpha\left[\frac{p}{2}(c_{\uparrow\uparrow}-c_{\uparrow\downarrow})+(1-p)\left\{\boxed{c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}}-c_{\uparrow\downarrow}\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}\right\}\right]$
$\frac{dc_{\downarrow\uparrow}}{dt}$ $=(1-\alpha)\left[-pc_{\downarrow\uparrow}+(1-p)\left\{c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}-c_{\downarrow\uparrow}\left(1-\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right)\right\}\right]$
$+\alpha\left[\frac{p}{2}(c_{\downarrow\downarrow}-c_{\downarrow\uparrow})+(1-p)\left\{\boxed{c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}}-c_{\downarrow\uparrow}\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right\}\right]$
$\frac{dc_{\downarrow\downarrow}}{dt}$ $=(1-\alpha)\left[pc_{\uparrow\downarrow}+(1-p)\left\{c_{\uparrow\downarrow}\left(1-\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}\right)-c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}\right\}\right]$
$+\alpha\left[\frac{p}{2}(c_{\downarrow\uparrow}-c_{\downarrow\downarrow})+(1-p)\left\{c_{\downarrow\uparrow}\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\boxed{-c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}}\right\}\right]$ (27)
and thus, according to definitions (3) and (4), the evolution of aggregate concentrations of positive public and positive private opinions can be expressed as:
$\frac{dc_{S}}{dt}$ $=(1-\alpha)(1-p)\left[\left\{c_{\downarrow\uparrow}\left(1-\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right)+c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}-c_{\uparrow\downarrow}\left(1-\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}\right)-c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}\right\}\right]$
$+(1-\alpha)p\left(c_{\downarrow\uparrow}-c_{\uparrow\downarrow}\right),$
$\frac{dc_{\sigma}}{dt}$ $=\alpha(1-p)\left\{\boxed{c_{\downarrow\downarrow}\left(\theta_{\downarrow\downarrow}^{\uparrow\uparrow}+\theta_{\downarrow\downarrow}^{\uparrow\downarrow}\right)^{q}}+c_{\uparrow\downarrow}\left(\theta_{\uparrow\downarrow}^{\uparrow\uparrow}+\theta_{\uparrow\downarrow}^{\uparrow\downarrow}\right)^{q}-c_{\downarrow\uparrow}\left(\theta_{\downarrow\uparrow}^{\downarrow\uparrow}+\theta_{\downarrow\uparrow}^{\downarrow\downarrow}\right)^{q}\boxed{-c_{\uparrow\uparrow}\left(\theta_{\uparrow\uparrow}^{\downarrow\uparrow}+\theta_{\uparrow\uparrow}^{\downarrow\downarrow}\right)^{q}}\right\}$
$+\alpha\frac{p}{2}\left(c_{\downarrow\downarrow}+c_{\uparrow\downarrow}-c_{\downarrow\uparrow}-c_{\uparrow\uparrow}\right)$ (28)
where $\theta_{X}^{Y}$ denotes the probability that a randomly selected neighbor of a target agent in state $X$ is in state $Y$. In pair approximation, this probability is modeled via concentration of edges between all possible agent states:
$\theta_{X}^{Y}=\frac{e_{X}^{Y}}{e_{X}^{\uparrow\uparrow}+e_{X}^{\uparrow\downarrow}+e_{X}^{\downarrow\uparrow}+e_{X}^{\downarrow\downarrow}},$ (29)
with $e_{X}^{Y}:=E_{X}^{Y}/E$, where $E_{X}^{Y}$ is the number of directed edges starting at an agent in state $X$ and ending at an agent is state $Y$; and $E$ is the total number of edges in the system. Since we consider undirected networks, $e_{X}^{Y}=e_{Y}^{X}$ holds for any $X$ and $Y$, which requires us to track only 10 unique edge concentrations. Additionally, exploiting the fact that the sum of concentrations must equate to 1, a total of 9 independent edge concentrations have to be considered. The time evolution of edge concentrations are in general described by the following equation
$\frac{de_{X}^{Y}}{dt}=\frac{1}{\langle k\rangle}\sum_{Z\neq W}c_{Z}\sum_{\kappa}P(\kappa\mid Z)f^{Z\rightarrow W}(\kappa)\Delta E_{X}^{Y}\mid^{Z\rightarrow W}(\kappa),$ (30)
where $\langle k\rangle$ is the global average node degree, $\kappa=[k_{\uparrow\uparrow},k_{\uparrow\downarrow},k_{\downarrow\uparrow},k_{\downarrow\downarrow}]$} is a vector of state populations in the direct neighborhood, denoting the number of neighbors in each state; $P(\kappa\mid Z)$ is the probability that an agent in state $Z$ has neighborhood $\kappa$; $f^{Z\rightarrow W}(\kappa)$ is the probability of an agent changing from state $Z$ to $W$ conditional on its neighborhood $\kappa$; and $\Delta E_{X}^{Y}\mid^{Z\rightarrow W}(\kappa)$ is the elementary change in the number of edges from agents in state $X$ to agents in $Y$, which occurs when an agent changes from $Z$ to $W$. The latter is given as
$\Delta E_{X}^{Y}\mid^{Z\rightarrow W}(\kappa)=k_{Y}(\delta_{W,X}-\delta_{Z,X})+k_{X}(\delta_{W,Y}-\delta_{Z,Y}),$ (31)
where $\delta_{i,j}$ is the Kronecker delta.
Let us explicitly list the flipping probabilities for all possible transitions between agent states. Most of these probabilities are identical in the models with and without self-anticonformity and will be presented later. The only differences occur for the transitions $\uparrow\uparrow\rightarrow\uparrow\downarrow$ and $\downarrow\downarrow\rightarrow\downarrow\uparrow$, so we start with these cases.
For the model with self-anticonformity:
$f^{\uparrow\uparrow\rightarrow\uparrow\downarrow}\left(\kappa\right)$ $=\alpha\left(1-p\right)\binom{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}}{q}\frac{1}{\binom{k}{q}}\mathbf{1}_{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\geqslant q}+\frac{1}{2}\alpha p$
$=\alpha\left(1-p\right)\frac{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\geqslant q}$
$+\frac{1}{2}\alpha p,$ (32)
$f^{\downarrow\downarrow\rightarrow\downarrow\uparrow}\left(\kappa\right)$ $=\alpha\left(1-p\right)\frac{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\geqslant q}$
$+\frac{1}{2}\alpha p,$ (33)
while without self-anticonformity we get:
$f^{\uparrow\uparrow\rightarrow\uparrow\downarrow}\left(\kappa\right)$ $=\frac{1}{2}\alpha p,$ (34)
$f^{\downarrow\downarrow\rightarrow\downarrow\uparrow}\left(\kappa\right)$ $=\frac{1}{2}\alpha p.$ (35)
The flipping probabilities of the remaining transitions are the same for both models:
$f^{\uparrow\uparrow\rightarrow\downarrow\uparrow}\left(\kappa\right)$ $=\left(1-\alpha\right)\left(1-p\right)\frac{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\geqslant q},$ (36)
$f^{\uparrow\downarrow\rightarrow\uparrow\uparrow}\left(\kappa\right)$ $=\alpha\left(1-p\right)\frac{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\geqslant q}$
$+\frac{1}{2}\alpha p,$ (37)
$f^{\uparrow\downarrow\rightarrow\downarrow\downarrow}\left(\kappa\right)$ $=\left(1-\alpha\right)$
$\times\left[1-\left(1-p\right)\frac{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\geqslant q}\right],$ (38)
$f^{\downarrow\uparrow\rightarrow\downarrow\downarrow}\left(\kappa\right)$ $=\alpha\left(1-p\right)\frac{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\geqslant q}$
$+\frac{1}{2}\alpha p,$ (39)
$f^{\downarrow\uparrow\rightarrow\uparrow\uparrow}\left(\kappa\right)$ $=\left(1-\alpha\right)$
$\times\left[1-\left(1-p\right)\frac{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\downarrow\uparrow}+k_{\downarrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\downarrow\uparrow}+k_{\downarrow\downarrow}\geqslant q}\right],$ (40)
$f^{\downarrow\downarrow\rightarrow\uparrow\downarrow}\left(\kappa\right)$ $=\left(1-\alpha\right)\left(1-p\right)\frac{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\right)!\left(k-q\right)!}{\left(k_{\uparrow\uparrow}+k_{\uparrow\downarrow}-q\right)!k!}\mathbf{1}_{k_{\uparrow\uparrow}+k_{\uparrow\downarrow}\geqslant q}.$ (41)
Finally, substituting Eqs. (31)-(41) into Eq. (30) allows us to obtain a system of differential equations that does not depend on the node degree distribution, but only on the average degree $\langle k\rangle_{X}$. The average node degree can be inferred from edge and state concentrations according to:
$\langle k\rangle_{X}=\frac{\sum_{Y}e_{X}^{Y}}{c_{X}}\langle k\rangle.$ (42)
As in other homogeneous pair-approximation treatments of $q$-voter models [14, 23], correlations between node degree and state do not emerge unless they are present in the initial condition. Hence, we can set $\langle k\rangle_{X}=\langle k\rangle$ for all $X$ and extract the opinion concentration from Eq. (42) according to:
$c_{X}=\sum_{Y}e_{X}^{Y},$ (43)
allowing us to fully describe the evolution of the system with $9$ edge concentrations. Detailed formulas for the differential equations describing the time evolution of the concentrations of all edges are included in the source codes, publicly available at github.com/BarbaraKaminska/EPO-q-voter.
IV Results
IV.1 Mean-field results
We begin by discussing the results for a complete graph, where all agents are mutually connected and can interact with each other. All-to-all interactions are consistent with the homogeneity assumed in the mean-field approximation, so in this setting the analytical predictions are expected to coincide with Monte Carlo simulations. We focus on the stationary values of three macroscopic quantities: the fraction of agents with positive expressed opinion, $c_{S}$, the fraction of agents with positive private opinion, $c_{\sigma}$, and the fraction of agents in dissonance, $d$ (see Eqs. (3)-(5)). We analyze these quantities as functions of the independence probability $p$ for the two versions of the $\alpha$-EPO model: with and without self-anticonformity.
As shown in Fig. 3, both models exhibit a phase transition between an ordered phase (agreement), in which the majority of agents share the same opinion ($c_{S}\neq 0.5$, $c_{\sigma}\neq 0.5$), and a disordered phase (disagreement), in which opinions are evenly split ($c_{S}=0.5$, $c_{\sigma}=0.5$). The location of the critical point $p^{*}$ and the type of the transition differ between the two $\alpha$-EPO variants. Nevertheless, within each variant, $c_{S}(p)$ and $c_{\sigma}(p)$ undergo the transition at the same critical point, and the transition is of the same type for both quantities. In the ordered phase ($p<p^{*}$), the majority opinion is more pronounced at the expressed level than at the private level. This result is consistent with pluralistic ignorance, i.e., a mismatch in which privately held views are more diverse than publicly expressed ones because individuals misperceive the prevailing norm and adjust their public behavior accordingly [27].
The role of $\alpha$ differs qualitatively between the two variants of the $\alpha$-EPO model. In the version with self-anticonformity, the stationary phase diagrams $c_{S}(p)$ and $c_{\sigma}(p)$ are independent of $\alpha$ and therefore coincide with those reported in Jędrzejewski et al. [13] (top row in Fig. 3). This is consistent with Eqs. (20)–(21), which, for $0<\alpha<1$, do not contain $\alpha$. In this variant, $\alpha$ affects only the stationary dissonance $d$, which increases with $\alpha$. This agrees with Jędrzejewski et al. [13], where dissonance was also the only quantity that distinguished the two update schemes: act then think (AT), in which the expressed opinion was updated first and the private opinion second, and think then act (TA), in which the private opinion was updated first and the expressed opinion second.
This robustness with respect to $\alpha$ is lost in the model without self-anticonformity (bottom row in Fig. 3). In this variant, the stationary curves $c_{S}(p)$, $c_{\sigma}(p)$, and $d(p)$ vary with $\alpha$. An intuitive explanation follows from the fact that private conformity is conditional on the expressed state in this model: even if the $q$-panel is unanimous, a private update is allowed only when the unanimous opinion of the panel coincides with the target’s expressed opinion $S_{i}$. Hence, the expressed layer effectively acts as a filter that can block private conformity events. As a consequence, changing $\alpha$ affects not only the relative frequency of updates at the private and expressed levels, but also the effectiveness of private conformity. Varying $\alpha$ can modify not only the location of the critical point $p^{*}$ but also the type of phase transition. For $q\leq 2$ the transition remains continuous, and for $q\geq 4$ it remains discontinuous, as in Kamińska and Sznajd-Weron [16]. However, for $q=3$ we observe a qualitative change: there exists a critical value $0.8\leq\alpha^{*}\leq 0.9$ above which the discontinuous transition becomes continuous.
IV.2 Pair approximation results
We now extend the analysis beyond complete graphs and account for network topology using the pair approximation. Within this approach we solve both $\alpha$-EPO variants, with and without strategic self-anticonformity, and compare the resulting phase behavior with the mean-field predictions. Overall, the pair approximation confirms the qualitative picture obtained in the mean-field analysis. In particular, the average degree $k$ only weakly affects the location of the transition: it slightly shifts the critical point $p^{*}$ for continuous transitions, and in the discontinuous case it shifts the boundaries of the hysteresis region.
For discontinuous transitions, the hysteresis region is delimited by two spinodals, $p_{\mathrm{low}}$ and $p_{\mathrm{up}}$, which mark the limits of stability of the ordered and disordered solutions, respectively. Figure 4 shows how $p^{*}{\mathrm{low}}$ and $p^{*}_{\mathrm{up}}$ depend on $\alpha$ for different influence-group sizes $q$ and average degrees $k$.
IV.3 Results on empirical networks
To test the robustness of our findings beyond idealized random or fully connected graphs, we performed Monte Carlo simulations on empirical social networks. Interestingly, as we show below, the pair approximation describes the simulation results surprisingly well.
We use the organizational networks collected by Fire et al. [7]. Although we conducted simulations on all networks reported in [7], here we present results only for two of them: the small network S2, consisting of $N=320$ nodes with average degree $k=14.81$ and average clustering coefficient $C=0.49$, and the medium network M2 with $N=3862$ nodes, $k=45.22$, and $C=0.31$. Average node degrees of the two selected networks are the closest to Dunbar’s number describing sizes of social circles (5, 15, 50 and 150 people who are increasingly distant). Furthermore, structure of these networks lacks hubs (nodes of high degree, whose nodes are sparsely connected with each other), that violate pair approximation assumptions.
Since the fractions of agents with positive expressed and private opinions, $c_{S}$ and $c_{\sigma}$, behave qualitatively similarly, in Fig. 5 we present only the stationary values of $c_{S}(p)$. We compare Monte Carlo results obtained on empirical networks with simulations on synthetic Watts–Strogatz networks constructed to match the corresponding empirical network in the number of nodes, average degree, and clustering coefficient, as well as with analytical predictions from the pair approximation.
Results for the small networks (panels a and d in Fig. 5) are noisy due to finite-size fluctuations. Despite this, we surprisingly find that for the S2 network the simulations on the empirical graph agree better with the analytical predictions than those on its Watts–Strogatz counterpart. For the M2 network (panels b and e in Fig. 5), the results are, first, much smoother. Second, they agree better with the analytical calculations, since this network has a lower clustering coefficient and thus weaker short-range correlations between neighboring nodes. Finally, in the last column of Fig. 5 we compare the pair-approximation results with simulations on random graphs of size $N=10^{4}$ and average degree $k=15,50,150$. This network class best matches the pair-approximation assumption of weak correlations between neighbors.
Figure 3: Stationary concentrations of positive opinions at public level $c_{S}$ (left column), private level $c_{\sigma}$ (middle column) and dissonance $d$ (right column) as a function of the probability of independence $p$ obtained from mean field approximation with influence group of size $q=3$. Rows correspond to different variants of the model: with (upper row) and without (bottom row) self-anticonformity. Results for $\alpha\in\{0.1,0.2,\ldots,0.9\}$ are shown, and the line color becomes more intense with increasing $\alpha$, i.e., the darker the line, the larger $\alpha$. In the upper row, the results for all values of $\alpha$ collapse onto the same curves for $c_{S}$ and $c_{\sigma}$; therefore, only a single line is visible.
Figure 4: Dependence of lower $p^{*}_{low}$ (dotted lines) and upper $p^{*}_{up}$ (dashed lines) spinodals on $\alpha$ obtained from pair approximation. Results are shown for both variants of the EPO model, with (top row) and without (bottom row) self-anticonformity, obtained within the pair approximation for networks with average degree $k=15$ (left column), $k=50$ (middle column), and $k=150$ (right column).
Figure 5: Stationary concentration of positive expressed opinions $c_{S}$ as a function of the independence probability $p$ for $\alpha=0.1,q=3$. Monte Carlo results on empirical networks [7] are compared with simulations on Watts–Strogatz networks matched in size, average degree, and clustering coefficient (panels a, b, d, e). Panels (c, f) show simulations on random graphs of size $N=10^{4}$ with average degree $k=15,50,150$. Full markers correspond to ordered initial conditions, while empty markers correspond to random initial conditions. Simulations results are compared with the pair-approximation prediction. Results are shown for both model variants: with (top row) and without (bottom row) self-anticonformity.
V Conclusions
In this paper, we introduced and compared two versions of the $\alpha$-EPO model, extending the two $q$-voter EPO variants proposed in [13, 16]. These two variants differ in that one includes self-anticonformity whereas the other does not. A parallel analysis of the corresponding $\alpha$-EPO versions allows us to investigate the role of self-anticonformity, which has recently attracted significant attention [5, 22]. We modified the original models by allowing internal and public states to evolve at different rates, which is motivated by the fact that people’s private and expressed opinions, beliefs and attitudes, or intentions and behaviors do not necessarily change simultaneously or equally often.
Both $\alpha$-EPO variants (with and without anticonformity) display a transition between an ordered (agreement) phase and a disordered (disagreement) phase. Although the position of the critical point $p^{*}$ and the character of the transition depend on the variant, within each model the expressed and private order parameters, $c_{S}(p)$ and $c_{\sigma}(p)$, undergo the transition at the same $p^{*}$ and with the same transition type. Importantly, in the ordered phase ($p<p^{*}$) agreement is systematically stronger at the expressed than at the private level, indicating pluralistic ignorance: public consensus can overstate the true degree of private agreement because agents misperceive the prevailing norm and adjust their expressed opinions accordingly.
In the variant with self-anticonformity, we find that the model is robust with respect to $\alpha$ as long as $0<\alpha<1$. Specifically, the stationary fractions of agents with positive expressed and private opinions, $c_{S}(p)$ and $c_{\sigma}(p)$, are identical for all $\alpha$. The only quantity that increases with $\alpha$ is the fraction of agents in dissonance, $d$, i.e., agents whose expressed and private opinions disagree. Hence, from an individual perspective, it may be beneficial for private opinions to be less volatile since fewer agents experience cognitive dissonance. At the same time, from a collective perspective, the update frequency does not matter because the stationary macroscopic state remains unchanged.
Once self-anticonformity is excluded, this robustness disappears: the stationary dependencies $c_{S}(p)$ and $c_{\sigma}(p)$ change with $\alpha$. In this variant, our results approach those of the original model [16] in the limit $\alpha\to 0$, which can be understood intuitively from the update rules: when updates occur predominantly at the expressed level, the expressed layer relaxes faster, and private conformity events are more likely to be effective because they are permitted only when the target’s expressed opinion already matches the unanimous influence group. For influence-group sizes $q\geq 2$, the critical point $p^{*}$ increases with $\alpha$, meaning that for larger $\alpha$ it is easier to reach an agreement; a majority of agents choose the same option even for a higher probability of independence $p$. Moreover, as $\alpha$ increases, the hysteresis region associated with a discontinuous phase transition becomes narrower.
For $q=3$, this effect is particularly clear: as $\alpha$ increases, the hysteresis region not only narrows but disappears entirely at a critical value $\alpha^{*}$, indicating a change from a discontinuous to a continuous phase transition. This reduction of hysteresis weakens bistability and path dependence, making collective outcomes less sensitive to initial conditions and less prone to delayed switching when conditions change. Thus, from the perspective of the whole group, in contrast to the individual level, it is advantageous when private attitudes are more volatile than publicly expressed behavior. This scenario is also consistent with psychological findings suggesting that people may change their attitudes more frequently than their behavior [19].
We show that the above observations hold not only on the complete graph (all-to-all interactions) but also on sparser artificial and empirical networks. Based on these results, we draw implications that may be relevant for managers and policymakers. First, it may be beneficial to encourage some people to play the devil’s advocate [10] and openly disagree with the majority. This result is particularly important in light of recent experiments in behavioral economics showing that standing out from the group can be a strategic response in situations where individuals are evaluated [5]. Second, it may be useful to discourage overly frequent changes in publicly expressed behavior. If open disagreement is encouraged and publicly expressed behavior is less volatile, agreement may become easier to reach and delayed responses to dynamically changing situations may be reduced.
We are aware that the above conclusions should be treated with caution, as they are derived from a specific class of models. An important direction for future work is to extend this type of analysis to other $\alpha$-EPO models—an umbrella term for models in which the interplay between private and expressed opinions is controlled by the parameter $\alpha$, which tunes the relative update rates of the two layers—in order to assess the generality of our conclusions and to better understand the role of anticonformity.
Acknowledgements.
This work was supported by funds from the National Science Centre (NCN, Poland) through grant no. 2023/51/I/HS6/02269.
Data Availability Statement
The source codes in C++ and MATLAB, which allow for the reproduction of all results in this publication, are publicly available on https://github.com/BarbaraKaminska/EPO-q-voter. The empirical network data used in this study is sourced from https://data4goodlab.github.io/dataset.html.